OpenStudy (anonymous):

if 2 angles are complmentary and one of the angles measures 27 degrees, what does the other angle measure? 27 63 73 153

OpenStudy (notamathgenius):

Well you need to know what a complimentary angle is, do you?

OpenStudy (anonymous):

the same right?

OpenStudy (lstrasmann):

Complementary angles are angles that add up to 90 degrees @Notamathgenius

OpenStudy (anonymous):

Their sum will be 90 degrees.. If one is 27 then what will be the other??

OpenStudy (anonymous):

\[x + 27 = 90\] Find x here..

OpenStudy (anonymous):

63

OpenStudy (lstrasmann):

Correct. And just for reference, supplementary angles add up to 180 degrees.

OpenStudy (anonymous):

let x be the complimentary angle,then x=90-27=63